2012 2013 45 331 111 197 255 573 416 1000 1573 1616 2012 2013 Assets Cash A/R Inventory Total Net Fixed Assets Total Assets 50 1200 Liabilities Accounts Payable157 308 Notes Payable Total Long-Term Debt 500 750 Common Stock Retained Earnings 95125 150 125 307 433 00 400 1302 1708 2013 Sales Cost of Goods Sold Depreciation EBIT Interest Taxes Net Income 3168 2534.4 200 433.6 80 60 293.6 What is the Current Ratio for 2013 (to nearest 2 decimals)?1.) What is the Current Ratio for 2013 (to nearest 2 decimals)?

2.) What is the Quick Ratio (to nearest 2 decimals)?

3.) What is the Times Interest Earned ratio for 2013 (nearest 2 decimals)?

4.) What is the Debt to Equity ratio for 2012 (to nearest 2 decimals)?

5.) What is the Debt to Equity ratio for 2013 (to nearest 2 decimals)?

6.) What is the Profit Margin for 2013 ( to nearest 2 decimals)?

7.) What is the Total Asset Turnover (to nearest two decimals)?

8.) What is the Equity Multiplier for 2013 (to nearest 2 decimals)?

9.) What is the Return on Equity for 2013 (to nearest 2 decimals)?

10.) What is the Return on Assets for 2013 (to the nearest 2 decimals)?
Answer #1

1

Current ratio = current assets/current liablities
Current ratio = 416/433
Current ratio = 0.96

2

Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (416-255)/433
Quick or acid test ratio = 0.37

3

Times interest earned ratio = EBIT/Interest
TIE ratio = 433.6/80
TIE ratio = 5.42

4

D/E = (debt+notes +accounts payable)/(common stock +retained earnings)

=(307+500)/(400+95)

=

1.63
Please ask remaining parts seperately